# Composite Learning Backstepping Control With Guaranteed Exponential Stability and Robustness

### v3

@@ Appendix B The proof of Theorem 2
$V(\bm{w})=\bm{e}^{T}\bm{e}/2+(1+p)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}/2$ (37)
with $\bm{w}$ :=:= $\bm{e}^{T},\tilde{\bm{\theta}}^{T}]^{T}\in\mathbb{R}^{n+N}$, where $p\in\mathbb{R}^{+}$ is a constant. There exist constants $\lambda_{\rm a}$ :=:= $\min\{1/2,(1+p)/(2\lambda_{\max}(\Gamma))\}$ and $\lambda_{\rm b}$ :=:= $\min\{1/2,(1+p)/(2\lambda_{\min}(\Gamma))\}$ such that
@@
$\dot{V}=(\dot{\bm{e}}^{T}\bm{e}+\bm{e}^{T}\dot{\bm{e}})/2+(1+p)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\dot{\tilde{\bm{\theta}}}$.
@@
$\dot{V}=$ $\bm{e}^{T}(-K\bm{e}+\Phi^{T}\tilde{\bm{\theta}})-(1+p)\tilde{\bm{\theta}}^{T}(\Phi_{{\rm f}}\Phi_{{\rm f}}^{T}+\kappa Q(t,t_{\rm e}))\tilde{\bm{\theta}}$ (42)
@@
$\dot{V}=$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa(1+p)\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}$
$+\bm{e}^{T}(\Phi-\Phi_{\rm f})\tilde{\bm{\theta}}-\bm{e}^{T}\bm{e}/4$
$-\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}+\bm{e}^{T}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}-p\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$
where the second line of the above formula satisfies
@@
$\dot{V}\leq$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa(1+p)\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}$
@@
$\dot{V}(t)\leq$ $-k_{\rm c}\|\bm{e}\|^{2}/2-k_{\rm c}(1+p)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}/2$
$+k_{\rm c}(1+p)c_{\theta}^{2}/(2\lambda_{\min}(\Gamma))+(\delta c_{\theta})^{2}/(2k_{\rm c})$
$=$ $-k_{\rm c}V(t)+k_{\rm c}(1+p)c_{\theta}^{2}/(2\lambda_{\min}(\Gamma))+(\delta c_{\theta})^{2}/(2k_{\rm c})$
$=$ $-k_{\rm c}V(t)/2-k_{\rm c}(V(t)-\eta(k_{\rm c},\alpha,\Gamma))/2$
with $\eta(k_{\rm c},\alpha,\Gamma)$ :=:= $(1+p)c_{\theta}^{2}/\lambda_{\min}(\Gamma)+(\delta c_{\theta}/k_{\rm c})^{2}\in\mathbb{R}^{+}$. Let $\Omega_{r}$ :=:= $\{{\bm{y}}_{{\rm r}n}|\dot{y}_{\rm r}$, $\cdots$, $y^{(n-1)}_{\rm r}\in\Omega_{{\rm c}_{\rm r}}\}\subset\mathbb{R}^{n}$, $\Omega_{{\rm c}_{w}}$ :=:= $\Omega_{{\rm c}_{x}}\cap\Omega_{r}\times\Omega_{{\rm c}_{\theta}}$, and $\Omega_{{\rm c}_{w0}}$ :=:= $\Omega_{{\rm c}_{0}}\cap\Omega_{r}\times\Omega_{{\rm c}_{\theta}}$ such that $\Omega_{{\rm c}_{w0}}\subset\Omega_{{\rm c}_{w}}$. It is implied from the above inequality that
@@
on $\bm{w}(t)\in\Omega_{{\rm c}_{w}}$ and $t\in[0,t_{\rm f})$. Based on the results presented in (38) and (45), the UUB Theorem [90, Th. 4.5] is applied to conclude that if $\eta<\sqrt{c_{w0}/\lambda_{\rm b}}$, then the closed-loop system (41) achieves UUB stability, where $\bm{e}(t)$, $\tilde{\bm{\theta}}(t)\in$ $L_{\infty}$, $\forall t\geq 0$ implying $t_{\rm f}=\infty$ so that the solution $\bm{w}(t)$ is unique, $\forall t\geq 0$ [90, Lemma 3.1]. It is clear from the definition of $\eta$ that given any $c_{0}\in\mathbb{R}^{+}$, there certainly exist suitable $k_{\rm c}$, $\alpha$, and $\Gamma$ such that $\eta<\sqrt{c_{w0}/\lambda_{\rm b}}$ holds. Besides, the transient bound for $\bm{e}$ derived by [90, Th. 4.5] is given as follows:
@@
Using Assumptions 1–2 and $\bm{e}(t)$, $\tilde{\bm{\theta}}(t)\in$ $L_{\infty}$, $\forall t\geq 0$, one gets $\bm{x}(t)$ $\in$ $L_{\infty}$, $\forall t\geq 0$ in (7). As the signals $\Phi_{\rm f}$, $\bm{z}$, $Q$, and $\bm{q}_{\rm f}$ in (23) are bounded, and are filtered by $H(s)$ in (16), their time derivatives up to the (n2n-2)th order are also bounded according to the input-to-state stability property of the stable filter $H(s)$ [87, 43]. As the high-order time derivatives of $\hat{\bm{\theta}}$ in (24) are computed as a weighted sum of these signals and their high-order time derivatives, one concludes ${\hat{\bm{\theta}}}^{(k)}\in L_{\infty}$ for all relevant $k$. Thus, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) is UUB stable on $t\in[0,\infty)$.
@@
$\left\{\begin{array}[]{l}\dot{\bm{e}}=\Lambda\bm{e}+\Phi_{\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}\\ \dot{\tilde{\bm{\theta}}}_{\zeta}=-\Gamma_{\zeta}\left(\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+\kappa Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}\right)\\ \dot{\tilde{\bm{\theta}}}_{0}=\bm{0}\in\mathbb{R}^{N-N_{\zeta}}\end{array}\right$. (50)
@@
$V_{\zeta}=\frac{1}{2}\bm{e}^{T}\bm{e}+\frac{1}{2}(1+p)\tilde{\bm{\theta}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}\tilde{\bm{\theta}}_{\zeta}$
@@
$\dot{V}_{\zeta}=(\dot{\bm{e}}^{T}\bm{e}+\bm{e}^{T}\dot{\bm{e}})/2+(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\Gamma^{-1}_{\zeta}\dot{\tilde{\bm{\theta}}}_{\zeta}.$
@@
$-(1+p)({\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+\kappa\tilde{\bm{\theta}}^{T}_{\zeta}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}).$ (51)
@@
$\dot{V}\!\leq\!-k_{\rm c}\bm{e}^{T}\bm{e}\!-\!\kappa(1\!+\!p)\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}\!+\!\bm{e}^{T}(\Phi_{\zeta}\!-\!\Phi_{{\rm f},\zeta})\tilde{\bm{\theta}}_{\zeta}.$ (52)
@@
$\dot{V}_{\zeta}\leq-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\bm{e}^{T}(\Phi_{\zeta}-\Phi_{\rm f,\zeta})^{T}\tilde{\bm{\theta}}_{\zeta}.$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|.$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta^{2}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}/(2k_{\rm c}).$
Choosing $p$ == $\delta^{2}/(2k_{\rm c}\kappa\sigma^{*})\in\mathbb{R}^{+}$, one obtains
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}$
$\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}/2-\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})\tilde{\bm{\theta}}^{T}_{\zeta}\Gamma_{\zeta}^{-1}\tilde{\bm{\theta}}_{\zeta}$
@@
with $k_{\rm p}$ :=:= $\min\{k_{\rm c},2\kappa\sigma^{*}\lambda_{\min}(\Gamma_{\zeta})/(1+p)\}\in\mathbb{R}^{+}$. Therefore, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) is partially exponentially stable on $t$ $\in$ $T_{\rm a},\infty)$, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to $\bm{0}$.
@@
$\dot{V}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa\sigma^{*}\lambda_{\min}(\Gamma)\tilde{\bm{\theta}}^{T}\Gamma^{-1}\tilde{\bm{\theta}}$
$\leq$ $-k_{\rm e}V,t\geq T_{\rm e}$
with $k_{\rm e}$ :=:= $\min\{k_{\rm c},2\kappa\sigma^{*}\lambda_{\min}(\Gamma)/(1+p)\}\in\mathbb{R}^{+}$ and $p$ == $\delta^{2}/$ $(2k_{\rm c}\kappa\sigma^{*})$ $\in\mathbb{R}^{+}$, where $\delta\in\mathbb{R}^{+}$ satisfies $\|\Phi-\Phi_{\rm f}\|$ $\leq$ $\delta$, $\forall t\geq 0$, and $\sigma^{*}\in\mathbb{R}^{+}$ with $\sigma^{*}\leq\sigma$ satisfies (30). Thus, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) is exponential stable on $t$ $\in$ $T_{\rm e},\infty)$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to $\bm{0}$. ∎

### v4

@@ Appendix B The proof of Theorem 2
$V(\bm{w})=\bm{e}^{T}\bm{e}/2+(1+p)\tilde{\bm{\theta}}^{T}\tilde{\bm{\theta}}/2$ (37)
with $\bm{w}$ :=:= $\bm{e}^{T},\tilde{\bm{\theta}}^{T}]^{T}\in\mathbb{R}^{n+N}$ and $p\in\mathbb{R}^{+}$ being a constant. There exist constants $\lambda_{\rm a}$ :=:= $1/2$ and $\lambda_{\rm b}$ :=:= $(1+p)/2$ to get
@@
$\dot{V}=(\dot{\bm{e}}^{T}\bm{e}+\bm{e}^{T}\dot{\bm{e}})/2+(1+p)\tilde{\bm{\theta}}^{T}\dot{\tilde{\bm{\theta}}}$.
@@
$\dot{V}=$ $\bm{e}^{T}(-K\bm{e}+\Phi^{T}\tilde{\bm{\theta}})-(1+p)\tilde{\bm{\theta}}^{T}(\Phi_{{\rm f}}\Phi_{{\rm f}}^{T}+\kappa_{2}Q(t,t_{\rm e}))\tilde{\bm{\theta}}$ (42)
@@
$\dot{V}=$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa_{2}(1+p)\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}$
$+\bm{e}^{T}(\Phi-\Phi_{\rm f})\tilde{\bm{\theta}}-p\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$
$-\bm{e}^{T}\bm{e}/4-\tilde{\bm{\theta}}^{T}\Phi_{\rm f}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}+\bm{e}^{T}\Phi_{\rm f}^{T}\tilde{\bm{\theta}}$
where the third line of the above formula satisfies
@@
$\dot{V}\leq$ $-\bm{e}^{T}(K-I/4)\bm{e}-\kappa_{2}(1+p)\tilde{\bm{\theta}}^{T}Q(t,t_{\rm e}){\tilde{\bm{\theta}}}$
@@
$\dot{V}(t)\leq$ $-k_{\rm c}\|\bm{e}\|^{2}/2-k_{\rm c}(1+p)\tilde{\bm{\theta}}^{T}\tilde{\bm{\theta}}/2$
$+k_{\rm c}(1+p)c_{\theta}^{2}/2+(\delta c_{\theta})^{2}/(2k_{\rm c})$
$=$ $-k_{\rm c}V(t)+k_{\rm c}(1+p)c_{\theta}^{2}/2+(\delta c_{\theta})^{2}/(2k_{\rm c})$
$=$ $-k_{\rm c}V(t)/2-k_{\rm c}(V(t)-\eta(k_{\rm c},\alpha))/2$
with $\eta(k_{\rm c},\alpha)$ :=:= $(1+p)c_{\theta}^{2}+(\delta c_{\theta}/k_{\rm c})^{2}\in\mathbb{R}^{+}$. Define sets $\Omega_{r}$ :=:= $\{{\bm{y}}_{{\rm r}n}|\dot{y}_{\rm r}$, $\cdots$, $y^{(n-1)}_{\rm r}\in\Omega_{{\rm c}_{\rm r}}\}\subset\mathbb{R}^{n}$, $\Omega_{{\rm c}_{w}}$ :=:= $\Omega_{{\rm c}_{x}}\cap\Omega_{r}\times\Omega_{{\rm c}_{\theta}}$, and $\Omega_{{\rm c}_{w0}}$ :=:= $\Omega_{{\rm c}_{0}}\cap\Omega_{r}\times\Omega_{{\rm c}_{\theta}}$ such that $\Omega_{{\rm c}_{w0}}\subset\Omega_{{\rm c}_{w}}$. It is implied from the above inequality that
@@
on $\bm{w}(t)\in\Omega_{{\rm c}_{w}}$ and $t\in[0,t_{\rm f})$. Based on the results presented in (38) and (45), the UUB Theorem [43, Th. 4.5] is applied to conclude that if $\eta<\sqrt{c_{w0}/\lambda_{\rm b}}$, then the closed-loop system (41) achieves UUB stability, where $\bm{e}(t)$, $\tilde{\bm{\theta}}(t)\in$ $L_{\infty}$, $\forall t\geq 0$ implying $t_{\rm f}=\infty$ so that the solution $\bm{w}(t)$ is unique, $\forall t\geq 0$ [43, Lemma 3.1]. It is clear from the definition of $\eta$ that given any $c_{0}\in\mathbb{R}^{+}$, there certainly exist suitable $k_{\rm c}$ and $\alpha$ such that $\eta<\sqrt{c_{w0}/\lambda_{\rm b}}$ holds. Besides, the transient bound for $\bm{e}$ derived by [43, Th. 4.5] is given as follows:
@@
Using Assumptions 1–2 and $\bm{e}(t)$, $\tilde{\bm{\theta}}(t)\in$ $L_{\infty}$, $\forall t\geq 0$, one gets $\bm{x}(t)$ $\in$ $L_{\infty}$, $\forall t\geq 0$ in (7). As the signals $\Phi_{\rm f}$, $\bm{z}$, $Q$, and $\bm{q}_{\rm f}$ in (23) are bounded, and are filtered by $H(s)$ in (16), their time derivatives up to the (n2n-2)th order are also bounded according to the input-to-state stability property of the stable filter $H(s)$ [40, 32]. As the high-order time derivatives of $\hat{\bm{\theta}}$ in (24) are computed as a weighted sum of these signals and their high-order time derivatives, one concludes ${\hat{\bm{\theta}}}^{(k)}\in L_{\infty}$ for all relevant $k$. Thus, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) has UUB stability on $t\in[0,\infty)$.
@@
$\left\{\begin{array}[]{l}\dot{\bm{e}}=\Lambda\bm{e}+\Phi_{\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}\\ \dot{\tilde{\bm{\theta}}}_{\zeta}=-\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}-\kappa_{2}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}\\ \dot{\tilde{\bm{\theta}}}_{0}=\bm{0}\in\mathbb{R}^{N-N_{\zeta}}\end{array}\right$. (50)
@@
$V_{\zeta}=\frac{1}{2}\bm{e}^{T}\bm{e}+\frac{1}{2}(1+p)\tilde{\bm{\theta}}_{\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}$
@@
$\dot{V}_{\zeta}=(\dot{\bm{e}}^{T}\bm{e}+\bm{e}^{T}\dot{\bm{e}})/2+(1+p){\tilde{\bm{\theta}}}_{\zeta}^{T}\dot{\tilde{\bm{\theta}}}_{\zeta}.$
@@
$-(1+p)({\tilde{\bm{\theta}}}_{\zeta}^{T}\Phi_{{\rm f},\zeta}\Phi_{{\rm f},\zeta}^{T}\tilde{\bm{\theta}}_{\zeta}+\kappa_{2}\tilde{\bm{\theta}}^{T}_{\zeta}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}).$ (51)
@@
$\dot{V}\!\leq\!-k_{\rm c}\bm{e}^{T}\bm{e}\!-\!\kappa_{2}(1\!+\!p)\tilde{\bm{\theta}}_{\zeta}^{T}Q_{\zeta}(t,t_{\rm e}){\tilde{\bm{\theta}}}_{\zeta}\!+\!\bm{e}^{T}(\Phi_{\zeta}\!-\!\Phi_{{\rm f},\zeta})\tilde{\bm{\theta}}_{\zeta}.$ (52)
@@
$\dot{V}_{\zeta}\leq-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa_{2}\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\bm{e}^{T}(\Phi_{\zeta}-\Phi_{\rm f,\zeta})^{T}\tilde{\bm{\theta}}_{\zeta}.$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa_{2}\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta\|\bm{e}\|\|\tilde{\bm{\theta}}_{\zeta}\|.$
@@
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}/2-\kappa_{2}\sigma^{*}(1+p)\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}+\delta^{2}\|\tilde{\bm{\theta}}_{\zeta}\|^{2}/(2k_{\rm c}).$
Choosing $p$ == $\delta^{2}/(2k_{\rm c}\kappa_{2}\sigma^{*})\in\mathbb{R}^{+}$, one obtains
$\dot{V}_{\zeta}\leq$ $-k_{\rm c}\bm{e}^{T}\bm{e}/2-\kappa_{2}\sigma^{*}\tilde{\bm{\theta}}^{T}_{\zeta}\tilde{\bm{\theta}}_{\zeta}$
@@
with $k_{\rm p}$ :=:= $\min\{k_{\rm c},2\kappa_{2}\sigma^{*}/(1+p)\}\in\mathbb{R}^{+}$, which implies that the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) achieves partial exponential stability on $t$ $\in$ $T_{\rm a},\infty)$, where the tracking error $\bm{e}$ and the partial estimation error $\tilde{\bm{\theta}}_{\zeta}$ exponentially converge to $\bm{0}$.
@@
$\dot{V}\leq-k_{\rm c}\bm{e}^{T}\bm{e}-\kappa_{2}\sigma^{*}\tilde{\bm{\theta}}^{T}\tilde{\bm{\theta}}\leq-k_{\rm e}V,\;t\geq T_{\rm e}$
with $k_{\rm e}$ :=:= $\min\{k_{\rm c},2\kappa_{2}\sigma^{*}/(1+p)\}\in\mathbb{R}^{+}$ and $p$ == $\delta^{2}/$ $(2k_{\rm c}\kappa_{2}\sigma^{*})$ $\in\mathbb{R}^{+}$, where $\delta\in\mathbb{R}^{+}$ satisfies $\|\Phi-\Phi_{\rm f}\|$ $\leq$ $\delta$, $\forall t\geq 0$, and $\sigma^{*}\in\mathbb{R}^{+}$ ($\sigma^{*}\leq\sigma$) satisfies (30). Hence, the equilibrium point $(\bm{e}$, $\tilde{\bm{\theta}})=\bm{0}$ of the closed-loop system (7) with (23) has exponential stability on $t$ $\in$ $T_{\rm e},\infty)$, where the tracking error $\bm{e}$ and the estimation error $\tilde{\bm{\theta}}$ exponentially converge to $\bm{0}$. ∎